OpenStudy (anonymous):

how do i start this problem? if f(x)=x^2+sinx, show that there must be a number c such that f(c)=50 help please

jimthompson5910 (jim_thompson5910):

is it f(c) = 50 OR is it f ' (c) = 50 ... notice the prime notation ??

OpenStudy (anonymous):

f(c)=50

jimthompson5910 (jim_thompson5910):

so it sounds like you need to solve f(x)=x^2+sinx f(c)=c^2+sin(c) 50 = c^2 + sin(c) 0 = c^2 + sin(c) - 50 you can't do this analytically, but you can get an approximation with a graphing calculator

jimthompson5910 (jim_thompson5910):

you don't need to actually solve for c though you just need to show that such a value of c exists to make 0 = c^2 + sin(c) - 50 true

jimthompson5910 (jim_thompson5910):

so that means you can use the intermediate value theorem do you know what I'm referring to?

OpenStudy (anonymous):

don't you need interval values to use ivt?

jimthompson5910 (jim_thompson5910):

yeah so you'll have to use a graphing calculator to spot the left and right boundaries of the interval

OpenStudy (anonymous):

i see im going to use geogebra be right back

jimthompson5910 (jim_thompson5910):

Use this if you don't have a graphing calculator https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

geogebra works too (actually works better)

jimthompson5910 (jim_thompson5910):

use x instead of c when it comes to graphing so graph x^2+sin(x)-50

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it must be from -7 to 7 i think

jimthompson5910 (jim_thompson5910):

you should see 2 x-intercepts

jimthompson5910 (jim_thompson5910):

what's a good interval for the x-intercept on the left

OpenStudy (anonymous):

yeah one at -7 and the other at 7

OpenStudy (anonymous):

on the left at -7

jimthompson5910 (jim_thompson5910):

it's close to -7, but not at -7

jimthompson5910 (jim_thompson5910):

the left x-intercept is between -8 and -7

OpenStudy (anonymous):

i see

jimthompson5910 (jim_thompson5910):

so if g(x) = x^2+sin(x)-50 then calculate g(-8) and g(-7) you should see a sign change

jimthompson5910 (jim_thompson5910):

the sign change will indicate there is at least one root for g(x) somewhere between -8 and -7

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

the other root is pretty close to +7, but not quite fully at +7 it's between +7 and +8 calculate g(7) and g(8) and you should see a sign change

OpenStudy (anonymous):

g(-8) is about 13.01 and g(-7) -1.66

jimthompson5910 (jim_thompson5910):

good and good

jimthompson5910 (jim_thompson5910):

the sign change from + to - indicates we have at least one root in between there

jimthompson5910 (jim_thompson5910):

assuming g(x) is continuous (which it is)

OpenStudy (anonymous):

g(7) is -0.34 and g(8) is 14.99

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so this shows that there are at least 2 values of c that make f(c)=50 true we have the graph to definitively say exactly 2, but you would use the mean value theorem to prove that (which would not involve the graph at all)

OpenStudy (anonymous):

ok